A, B, C and D are distinct digits, and 4 x AAB = CDA. If C is less than D, find the sum A+B+C+D.
Answer by ikleyn(52781)

.

Since AAB is 3-digit number and CDA is 3-digit number, there are only 2 possibilities for A:  A= 1  or  A= 2.


But A is the last digit of the number CDA, which is multiple of 4; so it leaves only one possibility for A to be 2:  A = 2.


Then  B = 3  or  B = 8.


Case 1).  If B = 3, then  4 * 223 = 892;  so,  C= 8  and  D= 9.

Case 2).  If B = 8, then  4 * 228 = 912;  so,  C= 9  and  D= 1.


The condition  C < D is satisfied in case 1 ONLY;  so, case 2) is denied.


Therefore,  A+B+c+D = 2+3+8+9 = 22.    ANSWER